The volume of an automobile tire is 2.5×10 −2 m3 . The pressure of the air in the tire is 2.5 atm and the temperature is 31C ◦ . what is is the mass of air in grams? The mean molecular mass of air is 29g
Answer #1

Solution : volume of tyre, V=2.5x10-2 m3 Peressure, P=2.5x101710=2.525x10 Pa Temperatore, T = 31+273= 304 K molecular mass of